# CoralVQA: A Large-Scale Visual Question Answering Dataset for Coral Reef Image Understanding

### v1

@@ 5.1.2 Coral VQA Tasks
We divide CoralVQA into three non-overlapping subsets to enable comprehensive evaluation: a training dataset, a testing dataset, and a cross-region dataset (marine region near Hawaii). Detailed statistics are presented in Table 2. We then establish three tasks related to coral reef conservation. (1) VQA on test dataset: Evaluation on the test dataset to establish standard performance benchmark. (2) VQA on cross-region dataset: Evaluation on the cross-region dataset, designed to examine model generalization to unseen regions.
@@
(3) VQA on bleaching-coverage dataset: Evaluation on our novel small-scale bleaching-coverage dataset (detailed in Section 5.4), which tests the model’s ability to perform a complex reasoning task. We fine-tune all the models on the training set and subsequently evaluate their performance on different tasks.
@@ 5.2 VQA Results on Test Dataset
Results As shown in Table 3, the following three conclusions can be drawn. First, compared to other methods, InternVL2.5 achieves superior average performance across both groups. Second, the accuracy of all models on open-ended questions is significantly lower than on closed-ended questions, with a performance gap exceeding 10%. For instance, the accuracy of InternVL2.5 for scene-type questions is 14.74% higher than for category-related questions. Similarly, Qwen2.5VL shows a 39.70% higher accuracy in addressing scene-type questions compared to category-related questions. These results suggest that open-ended questions often require domain-specific knowledge of coral categories and morphological characteristics, while current vision-language models pre-trained on general image-text datasets lack specialized knowledge in this area. Third, all methods show lower accuracy on questions regarding coral shape and quantity. In coral reef images, multiple coral genera are often distributed in interspersed patterns, with some genera being partially occluded. The limited ability of current visual-language models to extract textural, edge, and morphological features from coral reef images leads to enumeration errors, including both missed and duplicate counts.
@@
Methods Category Presence Quantity Color Position Size Shape Scene All Mini-Gemini 8.72 28.86 6.98 2.01 7.23 4.47 3.18 40.65 12.76 BLIP3 69.43 62.66 14.90 87.72 36.76 1.30 34.72 95.47 50.37 Qwen2.5VL 55.00 70.43 14.90 70.68 46.18 59.45 57.47 94.70 58.60 InternVL2.5 81.69 79.45 35.54 78.27 62.18 74.46 62.85 96.43 71.35 Methods Growth Algal Presence Quantity Integrity Susceptibility Environment Symbiosis All Mini-Gemini 38.74 27.88 30.61 20.96 37.26 38.71 24.82 53.34 34.04 BLIP3 88.09 61.67 94.51 63.47 96.54 80.92 74.83 82.60 80.32 Qwen2.5VL 78.52 58.18 90.56 60.50 72.56 74.37 72.00 72.45 72.39 InternVL2.5 86.96 65.45 88.73 65.25 95.90 80.04 80.20 88.81 81.42
@@
Methods Category Presence Quantity Color Position Size Shape Scene All Mini-Gemini 4.93 16.88 9.59 1.83 5.58 2.37 2.33 23.36 8.36 BLIP3 25.21 29.26 1.37 12.82 9.88 17.00 8.35 30.71 16.83 Qwen2.5VL 20.33 26.17 1.11 8.41 12.80 25.49 12.48 28.84 16.95 InternVL2.5 24.09 26.45 2.14 10.07 14.43 23.12 14.45 35.25 18.75 Methods Growth Algal Presence Quantity Integrity Susceptibility Environment Symbiosis All Mini-Gemini 18.65 15.48 18.84 17.13 17.38 - 11.78 20.64 17.13 BLIP3 32.35 26.03 32.06 8.80 31.07 - 38.03 33.80 28.88 Qwen2.5VL 31.82 24.10 30.12 2.78 26.04 - 33.50 31.54 25.70 InternVL2.5 31.93 26.21 31.20 6.25 29.89 - 36.58 35.71 28.25
@@ 5.3 VQA Results on Cross-Region Dataset
Settings and Evaluation Coral genera exhibit considerable variations in their composition, color, and morphology from different marine regions. Following the previous evaluation metrics, we adopt the accuracy as the metric to assess the model’s generalization performance on the cross-region dataset.
@@ 5.4 VQA Results on Bleaching-Coverage Dataset
Settings and Evaluation Evaluating the proportion of coral bleaching is essential for determining the extent of impacted ecosystems and the number of affected coral populations. Existing methods of coral bleaching coverage assessment typically rely on image segmentation to obtain the bleached coral area and calculate the coral bleaching coverage ratio. However, these methods heavily rely on domain knowledge and involve multiple steps, making them limited in application and time-consuming. In this context, VQA offers a promising alternative. The evaluation of coral bleaching rates through VQA presents significant challenges. This process requires not only pixel-level semantic segmentation and quantitative analysis of coral regions in images, but also a comprehensive understanding of both the spatial distribution and the visual details within coral images. In this section, we introduce a new VQA dataset called Bleaching-Coverage. Specifically, We collect 309 bleached coral reef images from the CoralVQA dataset and build the 309 question-answer pairs about coral bleaching coverage. We employ the LabelMe tool to segment the regions of bleached coral genera and calculate the proportion of the bleached area to the entire surveyed area, thereby obtaining the ground-truth. We utilize the MAE and MASE to evaluate the error between the output from VQA models and the ground truth.
Results As shown in Table 5, compared with Qwen2.5VL, InternVL2.5 has lower values of MAE and MASE, indicating better generalize ability for complex reasoning tasks. However, all methods still exhibit substantial prediction errors. The MASE for Qwen2.5VL and InternVL2.5 are 1.2326 and 0.8967, respectively. Furthermore, Mini-Gemini and BLIP3 fail to effectively comprehend the task-specific questions, often generating unrelated answers. These results suggest that existing visual-language models still face significant challenges in these complex tasks.
Table 5: Comparison of error metrics. The “-” denotes that the corresponding model can not generate valid answers.
Methods MAE MASE
Mini-Gemini - -
BLIP3 - -
Qwen2.5VL 0.1124 1.2326
InternVL2.5 0.0818 0.8967
@@ 5.5 Further Discussion
To further investigate the reason behind the model’s lower performance on questions related to coral size and shape, we conduct a case study. Focusing on InternVL2.5, we perform a visual analysis of its decision regions when answering questions about coral size and quantity. We design the following prompts to guide the model’s responses: “What is the genus of the largest hard coral in the image? Provide a brief response and output the location coordinates.” “How many coral genera are present in the image? Give a concise answer and specify their respective location coordinates.”
For coral size-related questions, the model incorrectly identifies non-target categories as the largest coral genus–mistakenly selecting regions labeled as other organisms, while the correct answer is the genus Porites, located at the top of the image. This mistake is mainly due to the model’s lack of domain-specific knowledge, which hinders its ability to distinguish between coral genera and similar-looking categories such as algae. For coral quantity-related questions, the model is unable to identify the positions of individual coral genera, leading to incorrect counts. These results highlight the model’s limitations in spatial reasoning, indicating that it lacks the fine-grained ecological understanding necessary for accurate differentiation among visually similar coral types. Furthermore, they demonstrate the necessity of developing novel frameworks for integrating marine biological knowledge into visual-language models to enhance their performance on coral reef analysis tasks.

### v2

@@ 5.1.2 Coral VQA Tasks
We divide CoralVQA into three non-overlapping subsets to enable comprehensive evaluation: a training dataset, a testing dataset, and a cross-region dataset (marine region near Hawaii). Detailed statistics are presented in Table 2. We then establish three tasks related to coral reef conservation. (1) VQA on test dataset: Evaluation on the test dataset to establish standard performance benchmark.
@@
(2) VQA on cross-region dataset: Evaluation on the cross-region dataset, designed to examine model generalization to unseen regions. (3) VQA on bleaching-coverage dataset: Evaluation on our novel small-scale bleaching-coverage dataset (detailed in Section 5.4), which tests the model’s ability to perform a complex reasoning task. We fine-tune all open-source models on the training set and subsequently evaluate their performance on different tasks.
@@ 5.2 VQA Results on Test Dataset
Results As shown in Table 3, the following three conclusions can be drawn. First, compared to other methods, InternVL2.5 achieves superior average performance across both groups. In addition, we further perform the experiments with two closed-source models: GPT-4o and Claude-3.5 Haiku. Although existing closed-source models have achieved good performance on general tasks, they perform poorly on coral-related tasks. Second, the accuracy of all models on open-ended questions is significantly lower than on closed-ended questions, with a performance gap exceeding 10%. For instance, the accuracy of InternVL2.5 for scene-type questions is 14.74% higher than for category-related questions. Similarly, Qwen2.5VL shows a 39.70% higher accuracy in addressing scene-type questions compared to category-related questions. These results suggest that open-ended questions often require domain-specific knowledge of coral categories and morphological characteristics, while current vision-language models pre-trained on general image-text datasets lack specialized knowledge in this area. Third, all methods show lower accuracy on questions regarding coral shape and quantity. In coral reef images, multiple coral genera are often distributed in interspersed patterns, with some genera being partially occluded. The limited ability of current visual-language models to extract textural, edge, and morphological features from coral reef images leads to enumeration errors, including both missed and duplicate counts.
@@
Methods Category Presence Quantity Color Position Size Shape Scene All GPT-4o 71.35 50.38 7.04 70.09 43.13 54.26 57.47 85.83 54.94 Claude3.5 Haiku 69.27 52.26 5.92 43.08 48.37 62.63 38.36 47.54 45.93 Mini-Gemini(FT) 8.72 28.86 6.98 2.01 7.23 4.47 3.18 40.65 12.76 BLIP3(FT) 69.43 62.66 14.90 87.72 36.76 1.30 34.72 95.47 50.37 Qwen2.5VL(FT) 55.00 70.43 14.90 70.68 46.18 59.45 57.47 94.70 58.60 InternVL2.5(FT) 81.69 79.45 35.54 78.27 62.18 74.46 62.85 96.43 71.35 Methods Growth Algal Presence Quantity Integrity Susceptibility Environment Symbiosis All GPT-4o 52.81 42.88 25.21 14.78 33.68 28.68 61.84 87.48 43.42 Claude3.5 Haiku 44.00 38.48 42.96 34.11 18.05 30.64 62.00 30.64 37.61 Mini-Gemini(FT) 38.74 27.88 30.61 20.96 37.26 38.71 24.82 53.34 34.04 BLIP3(FT) 88.09 61.67 94.51 63.47 96.54 80.92 74.83 82.60 80.32 Qwen2.5VL(FT) 78.52 58.18 90.56 60.50 72.56 74.37 72.00 72.45 72.39 InternVL2.5(FT) 86.96 65.45 88.73 65.25 95.90 80.04 80.20 88.81 81.42
@@
Methods Category Presence Quantity Color Position Size Shape Scene All Mini-Gemini(FT) 4.93 16.88 9.59 1.83 5.58 2.37 2.33 23.36 8.36 BLIP3(FT) 25.21 29.26 1.37 12.82 9.88 17.00 8.35 30.71 16.83 Qwen2.5VL(FT) 20.33 26.17 1.11 8.41 12.80 25.49 12.48 28.84 16.95 InternVL2.5(FT) 24.09 26.45 2.14 10.07 14.43 23.12 14.45 35.25 18.75 Methods Growth Algal Presence Quantity Integrity Susceptibility Environment Symbiosis All Mini-Gemini(FT) 18.65 15.48 18.84 17.13 17.38 - 11.78 20.64 17.13 BLIP3(FT) 32.35 26.03 32.06 8.80 31.07 - 38.03 33.80 28.88 Qwen2.5VL(FT) 31.82 24.10 30.12 2.78 26.04 - 33.50 31.54 25.70 InternVL2.5(FT) 31.93 26.21 31.20 6.25 29.89 - 36.58 35.71 28.25
@@ 5.3 VQA Results on Cross-Region Dataset
Settings and Evaluation Coral genera exhibit considerable intra-class variation in their composition, color, and morphology from different marine regions. Following the previous evaluation metrics, we adopt the accuracy as the metric to assess the model’s generalization performance on the cross-region dataset. Similar to the previous setting, we employ the GPT-4o API to assess the semantic consistency between cross-region dataset ground-truth answers and model-generated answers.
@@ 5.4 VQA Results on Bleaching-Coverage Dataset
Settings and Evaluation Evaluating the proportion of coral bleaching is essential for determining the extent of impacted ecosystems and the number of affected coral populations. Existing methods of coral bleaching coverage assessment typically rely on image segmentation to obtain the bleached coral area and calculate the coral bleaching coverage ratio. However, these methods heavily rely on domain knowledge and involve multiple steps, making them limited in application and time-consuming. In this context, VQA offers a promising alternative. The evaluation of coral bleaching rates through VQA presents significant challenges. This process requires not only pixel-level semantic segmentation and quantitative analysis of coral regions in images, but also a comprehensive understanding of both the spatial distribution and the visual details within coral images. In this section, we introduce a new VQA dataset called Bleaching-Coverage. Specifically, we collect 309 bleached coral reef images from the CoralVQA dataset and build the 309 question-answer pairs about coral bleaching coverage. We employ the LabelMe tool to segment the regions of bleached coral genera and calculate the proportion of the bleached area to the entire surveyed area, thereby obtaining the ground-truth. We utilize the MAE (Mean Absolute Error) and MASE (Mean Absolute Scaled Error) to evaluate the error between the output from VQA models and the ground truth.
Results The mean absolute error (MAE) of Qwen2.5VL is 0.1124, whereas that of InternVL2.5 is 0.0818. The MASE for Qwen2.5VL and InternVL2.5 are 1.2326 and 0.8967, respectively. Compared with Qwen2.5VL, InternVL2.5 has lower values of MAE and MASE, indicating better generalize ability for complex reasoning tasks. However, all methods still exhibit substantial prediction errors. Furthermore, Mini-Gemini and BLIP3 fail to effectively comprehend the task-specific questions, often generating unrelated answers. These results suggest that existing visual-language models still face significant challenges in these complex tasks. In addition, we visualize the bleaching regions of corals predicted by InternVL2.5 in the appendix, which further illustrates the limitations of existing vision-language models in assessing bleaching coverage.
@@ 5.5 Further Discussion
(a) Basic Visual Attributes.
(b) Ecological and Health-related Attributes
Figure 5: Performance comparison between fine-tuned and zero-shot evaluation on InternVL2.5 model. “All” represents the average accuracy across all question types.
Zero-shot Evaluation To further evaluate the real performance of existing visual-language models pre-trained on general image-text datasets in coral reef image understanding, we perform zero-shot evaluation on the InternVL2.5 models. Following the same evaluation method as before, we utilized the GPT-4o API to evaluate answer accuracy. As shown in Figure 5, two conclusions can be drawn. First, compared to the fine-tuning strategy, the average performance of zero-shot evaluation for InternVL2.5 shows varying degrees of degradation. For instance, InternVL2.5’s zero-shot evaluation shows performance decreases of 43.68 % and 35.54 % in the basic visual attributes, and ecological and health-related attributes, respectively. The result also indicated the effectiveness of fine-tuning strategies from the vision-language model. Second, for questions related to coral size, position, and quantity, the InternVL2.5’s question-answering accuracy falls below 20 %. The result indirectly reflects the limited capability of current vision-language models in extracting texture, edge, and morphological features from coral reef images. More detailed results are provided in the appendix. Case Study To further investigate the reason behind the model’s lower performance on questions related to coral size and shape, we conduct a case study. Focusing on InternVL2.5, as shown in Figure 6, we perform a visual analysis of its decision regions when answering questions about coral size and quantity. We design the following prompts to guide the model’s responses: “What is the genus of the largest hard coral in the image? Provide a brief response and output the location coordinates.” “How many coral genera are present in the image? Give a concise answer and specify their respective location coordinates.” For coral size-related questions, the InternVL2.5 model incorrectly identifies non-target categories as the largest coral genus–mistakenly selecting regions labeled as other organisms, while the correct answer is the genus Porites, located at the top of the image. This mistake is mainly due to the model’s lack of domain-specific knowledge, which hinders its ability to distinguish between coral genera and similar-looking categories such as algae. For coral quantity-related questions, the InternVL2.5 model is unable to identify the positions of individual coral genera, leading to incorrect counts. These results highlight the model’s limitations in spatial reasoning, indicating that it lacks the fine-grained ecological understanding necessary for accurate differentiation among visually similar coral types. Furthermore, they demonstrate the necessity of developing novel frameworks for integrating marine biological knowledge into visual-language models to enhance their performance on coral analysis tasks.
@@ 6 Conclusion and Future Works
7 Acknowledgments
This work is partially supported by National Natural Science Foundation of China under grants 62076232 and 62172049, BUPT Excellent Ph.D. Students Foundation under grant CX20251006, Science and Technology Development Foundation of South China Sea Bureau, Ministry of Natural Resources under grant 230208, Guangdong S&T Programme under grant 2025B1111130002, and Key Program of Marine Economy Development Special Foundation of Department of Natural Resources of Guangdong Province under grant GDNRC[2020]012. We sincerely thank the marine biologists in the fourth author’s research team (Nansha Islands Coral Reef Ecosystem National Observation and Research Station) for their help on data collection and processing.